# José Luis Blanco-Claraco

### v1

@@ 3.7 ICP optimization loop
  • Nearest neighbor search for each input point is generalized into a set of matching algorithms (see §3.7.1).
  • Several options are available as optimal pose solvers (see §3.7.2).
  • An optional metric map pipeline algorithm can be applied inside the ICP loop itself (see 3.7.3).
  • Dynamic variables (discussed in §3.6), including the ICP iteration count, can be used to parameterize any of the components above. This allows, for example, modifying the search radius over ICP iterations within one single time step (useful for loop closure).
@@
  • Importance is given to quality evaluators with a placeholder for such algorithms (see §3.7.4), fundamental in assessing whether a map registration is successful or not.
Keep in mind that the same ICP structure is to be used during online mapping (LiDAR odometry), localization, and for loop closure (see §3.13), although with different parameters and metric map layers.
@@
and $\varepsilon_{p}$ and $\varepsilon_{R}$ the corresponding thresholds, which are set to $10^{-4}$ and $5\cdot 10^{-5}$, respectively, in all presented experiments. For each iteration, all matching algorithms are sequentially invoked over the two input metric maps to populate a list of potential pairings $p$ (lines 11–14), then solvers are invoked using this list (lines 15–20) until a successful result $\mathbf{T}^{\prime}$ is obtained (see §3.7.2). The optional input hook for invoking a metric map updating pipeline is invoked next (line 21), which enables our claimed new feature of fluidly deskewing the input point cloud as the estimated pose is updated within the ICP iterations (see §3.7.3). The loop is run until convergence, then quality evaluation algorithms are applied (lines 29–35) and their result is returned together with the estimated pose.
@@ 3.7.1 Matchers
Regarding efficient parallelized implementation of matchers, we follow the findings reported in KISS-ICP (Vizzo et al. (2023)) and SIMPLE (Bhandari et al. (2024)) about the superiority of Intel Threading Building Blocks (TBB) (Reinders (2007)) over other alternatives to exploit multi-core CPUs, hence our implementation also employs that library to parallelize inner loops of most algorithms handling point clouds. A more detailed description of all available matchers and benchmarking them against different metric map data structures is out of the scope of this paper and is left for future works.
@@ 3.7.2 Solvers
: Having multiple solvers is in order due to the diversity of geometric entity combinations that can be found by matchers. In the following we are only focusing on point-to-point pairs since it has been shown to lead to good results (Vizzo et al. (2023)) although using diverse geometric primitives is worth further investigation, e.g. see MULLS (Pan et al. (2021)).
@@
is a classic problem with well-known closed-form solutions such as the quaternion method in Horn (1987). Unfortunately, in practice all pairings are not to be trusted equally due to the uneven and discrete nature of LiDAR sampling, sensor noise, and the existence of dynamic objects. Although our system offers Horn’s algorithm among the possible solvers, it is far from being the optimal choice Recent research has proven that robust (non least-squares) problem formulations lead to superior performance in the presence of pairing outliers and local minima, either using truncated least-squares and semidefinite programming (SDP) relaxation (Briales and Gonzalez-Jimenez (2017); Yang and Carlone (2019, 2020)) or by simply applying a robust kernel $\rho(\cdot)$ (e.g. Huber, Geman-McClure,…) to Eq. (6) (Chebrolu et al. (2021)):
$\boldsymbol{T}\hskip 2.29996pt^{\star}=\arg\min_{\boldsymbol{T}}\sum_{i=1}^{N}\rho(\left\lVert\mathbf{e}_{i}(\boldsymbol{T})\right\rVert^{2})$ (7)
In this work we follow this latter approach for its simplicity and superior efficiency in terms of running time. Furthermore, we propose adding an additional term to the cost function integrating an optional probability distribution reflecting any prior knowledge about the relative map poses. In the context of LO, this term incorporates vehicle pose predictions from the kinematic state estimator (block #2 in Figure 3, see §3.9), while in the context of loop closure, it reflects uncertainty built up along a given topological loop (see §3.13). In any case, such prior information is modeled as a Gaussian distribution with mean $\bar{\mathbf{T}}_{p}\in SE(3)$ and covariance matrix $\boldsymbol{\Sigma}_{p}$. This prior term, denoted as $\mathbf{e}_{p}$, is not affected by the robust kernel and is introduced as a Mahalanobis distance of the vee-operator applied to the Lie group logarithm map of the pose mismatch, that is:
@@
$c(\boldsymbol{T})$ $=$ ${\left\lVert\mathbf{e}_{p}(\boldsymbol{T})\right\rVert}_{\boldsymbol{\Sigma}_{p}}^{2}+\sum_{i=1}^{N}\rho(\left\lVert\mathbf{e}_{i}(\boldsymbol{T})\right\rVert^{2})$ (10)
@@
Solving Eq. (9)–(10) is a non-linear optimization problem, hence we propose using the iterative Gauss-Newton (GN) algorithm. Since the search space is non-Euclidean, the Lie group perturbation-model approach is employed (Sola et al. (2018); Blanco-Claraco (2021)) where the unknown becomes an increment ${\boldsymbol{\xi}}\in\mathfrak{se}(3)$ (SE(3) Lie group tangent Euclidean space):
@@ 3.7.3 Local map update pipeline
: As shown in line 21 of Algorithm 1, an optional metric map pipeline (§3.4) hook can be invoked within each ICP loop iteration. This feature is not used for loop-closure, but it is key for sequential processing of scans in odometry. In the default configuration for 3D LiDAR odometry (§3.8) this hook is used for two operations: (i) refining the estimated vehicle linear and angular velocities, and (ii) using such new estimated velocities to apply the de-skewing filter (§3.4.1) to the final layers of the metric map built from the current LiDAR scan (refer to “Observation pipeline (2/2)” in Figure 6).
@@
In practice, this simple method makes the velocity vectors to be estimated alongside with the vehicle incremental pose, enabling fluidly de-skewing points during ICP iterations so they better match the reference local map. As shown in the experimental results, this allows the effective tracking of abrupt motion profiles (see, e.g. §4.9).
Our approach has a similar aspiration than Continuous-Time ICP (Dellenbach et al. (2022)), but there are two differences: (i) ours does not introduce pose discontinuities between the end of a scan and the beginning of the next one, and (ii) ours does not need to use heuristic weights for the different parts of the optimization (point registration errors, final pose, and velocity vectors) since velocities are implicitly defined from the iteratively refined final pose, and the latter is unconstrained.
@@ 3.7.4 Quality estimators
: Except in degenerate cases, all solvers discussed above such as the classical quaternion method (Horn (1987)) or the GN iterative solver will always produce a result, correct or wrong depending on the existence of outliers in the input pairings. It hence becomes essential to assess whether the output relative pose between the two metric maps is plausible or should otherwise be discarded. In this latter case, our framework discards those key-frames with an unacceptable ICP quality. If this happens right at the beginning of mapping, the metric map is discarded and started from scratch again. Otherwise, amid navigation, the kinematic state estimator extrapolates the predicted vehicle motion to try to recover a good pose tracking in subsequent frames.
@@
Figure 5: Heuristic loss function to compare voxel occupancies $p_{i},p_{j}~{}\in[0,1]$ from two maps $i$ and $j$ in the proposed alignment quality metric in §3.7.4. The exact equation represented here is $L(p_{i},p_{j})=1.5+p_{i}+p_{j}-12p_{i}^{2}+22p_{i}p_{j}-12p_{j}^{2}$.
Apart of these quality metrics, note that the Gauss-Newton optimizer described in §3.7.2 also provides an uncertainty estimation for the solution, i.e. the Hessian matrix of the last iteration. Although normally overconfident, its eigenvalues and eigenvectors can be also used to assess whether a given relative pose is well-conditioned.
Figure 6: Detailed view of the “observation pipeline” blocks in Figure 3 for the “default 3D LiDAR” LO system configuration. Refer to discussion on pipelines in general in §3.4 and to the rationale behind this particular pipeline in §3.8.
@@ 3.8.2 Observation pipelines
: Split in two parts, (1/2) and (2/2), with contents detailed in Figure 6. Recall that the pipeline is split to allow the first part to be invoked only once per incoming scan, and the latter once for each ICP iteration. As sketched in the figure, a default generator first takes the raw sensor data and produces a point cloud map layer (“raw”) with contiguous memory layout (see §3.3). If per-point timestamps are present, they are then adjusted, with the particular convention of using the middle of the scan as the time origin. Next, the cloud is down-sampled and nearby points removed (FilterByRange in the figure) to avoid polluting the map with vehicle or robot body measurements. Next, a bounding box filter is applied to remove nearby ceiling points, which have been shown to lead to divergence in some scenarios. Then, this point cloud, which underwent decimation once, is decimated once again so we have two layers: a denser one to update the local map, and the sparser one to be used for registration in the ICP algorithm. This idea of using dual-resolution maps comes from its successful implementation in KISS-ICP (Vizzo et al. (2023)). Finally, note that de-skewing happens at the end, once for each down-sampled clouds, whereas all former works perform this step at the beginning instead. The reason for this alternative placement is twofold: to enable the fluid de-skewing discussed in §3.7.3, and increasing the efficiency since far less points are de-skewed in comparison to performing this operation at the very beginning of the pipeline. The result of first down-sampling then de-skewing is not exactly the same than switching the order of both operations, but in practice both methods lead to equivalent overall trajectory quality metrics.
@@ 3.8.3 Local map update pipeline
: This pipeline comprises just one map insertion action (§3.4.5), taking one of the observation map layer (the denser cloud) and merging it into the local metric map, effectively making the map to grow as the vehicle explores the environment.
@@ 3.8.4 ICP-related blocks
: One matcher algorithm is defined to find out point-to-point correspondences between the local map and one of the observation layers (the sparser cloud). Then, the Gauss-Newton solver described in §3.7.2 is defined as the unique solver. Following KISS-ICP (Vizzo et al. (2023)), the matcher threshold and the solver robust kernel are parameterized with a dynamic threshold, although estimated in a different way (see §3.6.3). Finally, quality for LO is assessed by using the matching ratio criterion (§3.7.4).
Figure 7: Example configuration for building metric maps from view-based maps (using the application sm2mm): a basic pipeline for building one single point cloud by accumulating decimated versions of de-skewed LiDAR scans. Refer to discussion in §3.11.
@@ 3.9 Kinematic state prediction
Figure 8: Example configuration for building metric maps from view-based maps (using the application sm2mm): a more advanced pipeline for classification of points into static or dynamic via volumetric probabilistic occupancy. See output examples in Figure 9 and discussion in §3.11.
3.10 “Simple-maps”: implementing view-based maps
This section provides some insights into how view-based maps (see §3.1) are implemented in our framework. As stated above, these maps, dubbed “simple-maps” in the framework, are the output of LO, hence they must include all the required information to build actual metric maps from them in the most flexible way. Thus, each such map $\mathfrak{m}$ is defined as a sequence of $N$ key-frames $\mathbf{k}_{i}$:
@@
Within observations we include an additional metadata observation which, instead of coming from a real sensor, is filled during LO. It includes information such as the bounding box, in robocentric coordinates, of all sensed points for each keyframe. This enables efficient determination of unfeasible loop-closures without having to analyze the raw sensor data (see §3.13). Furthermore, given that simple-maps may make use of the lazy-load feature described in §3.2 to make them efficient to load and parse, this metadata enables loading from disk only those sensor frames that actually are needed for loop-closure hypothesis checking.
@@
Figure 9: Layers of the metric map built using the pipeline of Figure 8 from the Luxembourg Garden automotive dataset in Dellenbach et al. (2022). Note how most traces of moving pedestrians and vehicles in (a) have been correctly removed in (b). See discussion in §3.11.
3.11 Simple-map conversion to metric maps
Once a simple-map is built, one of the most useful tasks to perform with them is constructing particular kinds of metric maps from it; this corresponds to the task represented in Figure 1(c) while discussing the framework overview.
@@ 3.11 Simple-map conversion to metric maps
The utility of such flexible design is illustrated with two example metric map pipelines [^4] Their corresponding YAML file descriptions can be checked out online under demos in the mp2p_icp project.. Possibly the simplest pipeline for 3D LiDAR datasets, sketched in Figure 7, uses a generator converting raw observations into a “raw” point cloud, which is then de-skewed after adjusting per-point timestamps to a given fixed convention. Recall de-skew in this context is possible since estimated velocities are stored within simple-maps (see §3.10). Next, scan points are down-sampled and spatially filtered to remove observations from the vehicle body. Finally, all layers are cleared so they start empty for the next map frame, except the map layer used as output, where all points accumulate.
A more elaborate pipeline is presented in Figure 8, with the goal of generating point clouds where mobile objects are removed, which is accomplished as follows. On the left, we have three generators: one to convert raw sensory data into point clouds (just as in the former pipeline), plus other two generators, each only invoked once to create two of the final metric maps. Note that each generator will create metric map layers of different types and with a particular sets of parameters (e.g. resolution, log-odds occupancy update weights, etc.). These two maps are a point cloud (layer “map”) and a 3D occupancy voxel map (§3.3). Raw points are de-skewed and spatially filtered to remove self-body interferences. That version of the cloud is inserted, for each frame, into the final point cloud “map”, whereas its down-sampled version (for efficiency reasons) is used to perform ray-tracing and update the occupancy value of the voxel map. Once all map frames are processed, the final filtering stage makes use of the filter described in §3.4.6 to split the map point cloud into those likely to be static and dynamic. As an example of the obtained results, Figure 9 shows the output map layers for a small fragment of the Paris LuCo dataset (Dellenbach et al. (2022)), where it can be seen how most 3D points corresponding to pedestrians, vehicles, or noisy points (e.g. near the edges of light poles), are removed from the static layer. Note that since there is no filter implementing learning-based semantic segmentation yet, vehicles and pedestrians that remain still are classified as static objects.

### v2

@@ 3.6 ICP optimization loop
  • Nearest neighbor search for each input point is generalized into a set of matching algorithms (see Section 3.6.1).
  • Several options are available as optimal pose solver algorithms from a set of predefined correspondences (see Section 3.6.2).
  • An optional metric map pipeline algorithm can be applied inside the ICP loop itself (see 3.6.3).
  • Dynamic variables (discussed in Section 3.5), including the ICP iteration count, can be used to parameterize any of the components above. This allows, for example, modifying the search radius over ICP iterations within one single time step (useful for loop closure).
@@
  • Importance is given to quality evaluators with a placeholder for such algorithms (see Section 3.6.4), fundamental in assessing whether a map registration is successful or not.
Keep in mind that the same ICP structure is to be used during online mapping (LiDAR odometry), localization, and for loop closure (see Section 3.12), although with different parameters and metric map layers.
@@
and $\varepsilon_{p}$ and $\varepsilon_{R}$ the corresponding thresholds, which are set to $10^{-4}$ and $5\cdot 10^{-5}$, respectively, in all presented experiments. For each iteration, all matching algorithms are sequentially invoked over the two input metric maps to populate a list of potential pairings $p$ (lines 11–14), then solvers are invoked using this list (lines 15–20) until a successful result $\mathbf{T}^{\prime}$ is obtained (see Section 3.6.2). The optional input hook for invoking a metric map updating pipeline is invoked next (line 21), which enables our claimed new feature of fluidly deskewing the input point cloud as the estimated pose is updated within the ICP iterations (see Section 3.6.3). The loop is run until convergence, then quality evaluation algorithms are applied (lines 29–35) and their result is returned together with the estimated pose.
@@ 3.6.1 Matchers
Regarding efficient parallelized implementation of matchers, we follow the findings reported in KISS-ICP (Vizzo et al., 2023) and SIMPLE (Bhandari et al., 2024) about the superiority of Intel Threading Building Blocks (TBB) (Reinders, 2007) over other alternatives to exploit multi-core CPUs, hence our implementation also employs that library to parallelize inner loops of most algorithms handling point clouds. A more detailed description of all available matchers and benchmarking them against different metric map data structures is out of the scope of this paper and is left for future works.
@@ 3.6.2 Solvers
: Having multiple solvers is in order due to the diversity of geometric entity combinations that can be found by matchers. In particular, our implementation offers the classic Horn’s method in Horn (1987) (apt for point-to-point pairings) and a generic Gauss-Newton solver suitable for different types of geometries. In the following we are focusing on point-to-point pairings since they have been shown to lead to good results (Vizzo et al., 2023). Point-to-plane correspondences are also used for our 3D-NDT system configuration, but their equations are left out for space reasons and will be described somewhere else. More diverse geometric primitives are worth further investigation, e.g. see MULLS (Pan et al., 2021).
@@
is a classic problem with well-known closed-form solutions such as the quaternion method in Horn (1987). Unfortunately, in practice all pairings are not to be trusted equally due to the uneven and discrete nature of LiDAR sampling, sensor noise, and the existence of dynamic objects. Although our system offers Horn’s algorithm among the possible solvers, it is far from being the optimal choice (see Section 7.4). Recent research has proven that robust (non least-squares) problem formulations lead to superior performance in the presence of pairing outliers and local minima, either using truncated least-squares and semidefinite programming (SDP) relaxation (Briales and Gonzalez-Jimenez, 2017; Yang and Carlone, 2019, 2020) or by simply applying a robust kernel $\rho(\cdot)$ (e.g. Huber, Geman-McClure,…) to Eq. (6) (Chebrolu et al., 2021):
$\boldsymbol{T}\hskip 2.29996pt^{\star}=\arg\min_{\boldsymbol{T}}\sum_{i=1}^{N}\rho(\left\lVert\mathbf{e}_{i}(\boldsymbol{T})\right\rVert)$ (7)
In this work we follow this latter approach for its simplicity and superior efficiency in terms of running time. Furthermore, we propose adding an additional term to the cost function integrating an optional probability distribution reflecting any prior knowledge about the relative map poses. In the context of LO, this term incorporates vehicle pose predictions from the kinematic state estimator (block #2 in Figure 3, see Section 3.8), while in the context of loop closure, it reflects uncertainty built up along a given topological loop (see Section 3.12). In any case, such prior information is modeled as a Gaussian distribution with mean $\bar{\mathbf{T}}_{p}\in SE(3)$ and covariance matrix $\boldsymbol{\Sigma}_{p}$. This prior term, denoted as $\mathbf{e}_{p}$, is not affected by the robust kernel and is introduced as a Mahalanobis distance of the vee-operator applied to the Lie group logarithm map of the pose mismatch, that is:
@@
$c(\boldsymbol{T})$ $=$ ${\left\lVert\mathbf{e}_{p}(\boldsymbol{T})\right\rVert}_{\boldsymbol{\Sigma}_{p}}^{2}+\sum_{i=1}^{N}\rho(\left\lVert\mathbf{e}_{i}(\boldsymbol{T})\right\rVert)$ (10)
@@
Solving Eq. (9)–(10) is a non-linear optimization problem, hence we propose using the iterative Gauss-Newton (GN) algorithm. Since the search space is non-Euclidean, the Lie group perturbation-model approach is employed (Sola et al., 2018; Blanco-Claraco, 2021) where the unknown becomes an increment ${\boldsymbol{\xi}}\in\mathfrak{se}(3)$ (SE(3) Lie group tangent Euclidean space):
@@ 3.6.3 Local map update pipeline
: As shown in line 21 of Algorithm 1, an optional metric map pipeline (Section 3.3) hook can be invoked within each ICP loop iteration. This feature is not used for loop-closure, but it is key for sequential processing of scans in odometry. In the default configuration for 3D LiDAR odometry (Section 3.7) this hook is used for two operations: (i) refining the estimated vehicle linear and angular velocities, and (ii) using such new estimated velocities to apply the de-skewing filter (Section 3.3.1) to the final layers of the metric map built from the current LiDAR scan (refer to “Observation pipeline (2/2)” in Figure 6).
@@
In practice, this simple method makes the velocity vectors to be estimated alongside with the vehicle incremental pose, enabling fluidly de-skewing points during ICP iterations so they better match the reference local map. As shown in the experimental results and in the ablation study in Section 7.3, this allows the effective tracking of abrupt motion profiles (see, e.g. Section 4.9).
Our approach has a similar aspiration than Continuous-Time ICP (Dellenbach et al., 2022), but there are two differences: (i) ours does not introduce pose discontinuities between the end of a scan and the beginning of the next one, and (ii) ours does not need to use heuristic weights for the different parts of the optimization (point registration errors, final pose, and velocity vectors) since velocities are implicitly defined from the iteratively refined final pose, and the latter is unconstrained.
@@ 3.6.4 Quality estimators
: Except in degenerate cases, all solvers discussed above such as the classical quaternion method (Horn, 1987) or the GN iterative solver will always produce a result, correct or wrong depending on the existence of outliers in the input pairings. It hence becomes essential to assess whether the output relative pose between the two metric maps is plausible or should otherwise be discarded. In this latter case, our framework discards those key-frames with an unacceptable ICP quality. If this happens right at the beginning of mapping, the metric map is discarded and started from scratch again. Otherwise, amid navigation, the kinematic state estimator extrapolates the predicted vehicle motion to try to recover a good pose tracking in subsequent frames.
@@
Figure 5: Heuristic loss function to compare voxel occupancies $p_{i},p_{j}~{}\in[0,1]$ from two maps $i$ and $j$ in the proposed alignment quality metric in Section 3.6.4. The exact equation represented here is $L(p_{i},p_{j})=1.5+p_{i}+p_{j}-12p_{i}^{2}+22p_{i}p_{j}-12p_{j}^{2}$.
Apart of these quality metrics, note that the Gauss-Newton optimizer described in Section 3.6.2 also provides an uncertainty estimation for the solution, i.e. the Hessian matrix of the last iteration. Although normally overconfident, its eigenvalues and eigenvectors can be also used to assess whether a given relative pose is well-conditioned.
Figure 6: Detailed view of the “observation pipeline” blocks in Figure 3 for the “default 3D LiDAR” LO system configuration. Refer to discussion on pipelines in general in Section 3.3 and to the rationale behind this particular pipeline in Section 3.7.
@@ 3.7.2 Observation pipelines
: Split in two parts, (1/2) and (2/2), with contents detailed in Figure 6. Recall that the pipeline is split to allow the first part to be invoked only once per incoming scan, and the latter once for each ICP iteration. As sketched in the figure, a default generator first takes the raw sensor data and produces a point cloud map layer (“raw”) with contiguous memory layout (see Section 3.2). If per-point timestamps are present, they are then adjusted, with the particular convention of using the middle of the scan as the time origin. Next, the cloud is down-sampled and nearby points removed (FilterByRange in the figure) to avoid polluting the map with vehicle or robot body measurements. Next, a bounding box filter is applied to remove nearby ceiling points, which have been shown to lead to divergence in some scenarios. Then, this point cloud, which underwent decimation once, is decimated once again so we have two layers: a denser one to update the local map, and the sparser one to be used for registration in the ICP algorithm. This idea of using dual-resolution maps comes from its successful implementation in KISS-ICP (Vizzo et al., 2023). Finally, note that de-skewing happens at the end, once for each down-sampled clouds, whereas all former works perform this step at the beginning instead. The reason for this alternative placement is twofold: to enable the fluid de-skewing discussed in Section 3.6.3, and increasing the efficiency since far less points are de-skewed in comparison to performing this operation at the very beginning of the pipeline. The result of first down-sampling then de-skewing is not exactly the same than switching the order of both operations, but in practice both methods lead to equivalent overall trajectory quality metrics.
@@ 3.7.3 Local map update pipeline
: This pipeline comprises just one map insertion action (Section 3.3.5), taking one of the observation map layer (the denser cloud) and merging it into the local metric map, effectively making the map to grow as the vehicle explores the environment.
@@ 3.7.4 ICP-related blocks
: One matcher algorithm is defined to find out point-to-point correspondences between the local map and one of the observation layers (the sparser cloud). Then, the Gauss-Newton solver described in Section 3.6.2 is defined as the unique solver. Following KISS-ICP (Vizzo et al., 2023), the matcher threshold and the solver robust kernel are parameterized with a dynamic threshold, although estimated in a different way (see Section 3.5.3). Finally, quality for LO is assessed by using the matching ratio criterion (Section 3.6.4).
Figure 7: Example configuration for building metric maps from view-based maps (using the application sm2mm): a basic pipeline for building one single point cloud by accumulating decimated versions of de-skewed LiDAR scans. Refer to discussion in Section 3.10.
@@ 3.8 Kinematic state prediction
Figure 8: Example configuration for building metric maps from view-based maps (using the application sm2mm): a more advanced pipeline for classification of points into static or dynamic via volumetric probabilistic occupancy. See output examples in Figure 9 and discussion in Section 3.10.
3.9 Implementation of view-based maps
This section provides some insights into how view-based maps (see Section 3.1) are implemented in our framework. As stated above, these maps are the output of LO, hence they must include all the required information to build actual metric maps from them in the most flexible way. Thus, each such map $\mathfrak{m}$ is defined as a sequence of $N$ key-frames $\mathbf{k}_{i}$:
@@ 3.9 Implementation of view-based maps
Within observations we include an additional metadata observation which, instead of coming from a real sensor, is filled during LO. It includes information such as the bounding box, in robocentric coordinates, of all sensed points for each keyframe. This enables efficient determination of unfeasible loop-closures without having to analyze the raw sensor data (see Section 3.12). Furthermore, given that these maps may make use of the lazy-load feature described in Section 3.13 to make them efficient to load and parse, this metadata enables loading from disk only those sensor frames that actually are needed for loop-closure hypothesis checking.
@@
Figure 9: Layers of the metric map built using the pipeline of Figure 8 from the Luxembourg Garden automotive dataset in Dellenbach et al. (2022). Note how most traces of moving pedestrians and vehicles in (a) have been correctly removed in (b). See discussion in Section 3.10.
PedestrianVehicle
(a) Layer: “map” (original map)
(b) Layer: “static”
(c) Layer: “dynamic”
Figure 10: Layers of the metric map built using the pipeline of Figure 8 from the sequence 00_4390_to_4530 in the ERASOR dataset (Lim et al., 2021). The estimated vehicle trajectory is shown as a thick black solid line. Note how traces of moving pedestrians and vehicles in (a) have been correctly removed in (b). The “dynamic” layer in (c) contains those moving objects, together with noisy points from environment voxels that were not consistently observed as occupied with a level of confidence high enough. See discussion in Section 3.10.
3.10 Converting view-based map to metric maps
Once a view-based map is populated with key-frames, one can use it to build a particular kind of metric map from it; this corresponds to the task represented in Figure 1(c) while discussing the framework overview.
@@ 3.10 Converting view-based map to metric maps
The utility of such flexible design is illustrated with two example metric map pipelines [^3] Their corresponding YAML file descriptions can be checked out online under demos in the mp2p_icp project.. Possibly the simplest pipeline for 3D LiDAR datasets, sketched in Figure 7, uses a generator converting raw observations into a “raw” point cloud, which is then de-skewed after adjusting per-point timestamps to a given fixed convention. Recall de-skew in this context is possible since estimated velocities are stored within map key-frames (see Section 3.9). Next, scan points are down-sampled and spatially filtered to remove observations from the vehicle body. Finally, all layers are cleared so they start empty for the next map frame, except the map layer used as output, where all points accumulate.
A more elaborate pipeline is presented in Figure 8, with the goal of generating point clouds where mobile objects are removed, which is accomplished as follows. On the left, we have three generators: one to convert raw sensory data into point clouds (just as in the former pipeline), plus other two generators, each only invoked once to create two of the final metric maps. Note that each generator will create metric map layers of different types and with a particular sets of parameters (e.g. resolution, log-odds occupancy update weights, etc.). These two maps are a point cloud (layer “map”) and a 3D occupancy voxel map (Section 3.2). Raw points are de-skewed and spatially filtered to remove self-body interferences. That version of the cloud is inserted, for each frame, into the final point cloud “map”, whereas its down-sampled version (for efficiency reasons) is used to perform ray-tracing and update the occupancy value of the voxel map. Once all map frames are processed, the final filtering stage makes use of the filter described in Section 3.3.6 to split the map point cloud into those likely to be static and dynamic. As a first example of the attainable results, Figure 9 shows the output map layers for a small fragment of the Paris LuCo dataset (Dellenbach et al., 2022), where it can be seen how most 3D points corresponding to pedestrians, vehicles, or noisy points (e.g. near the edges of light poles), are removed from the static layer. Note that since there is no filter implementing learning-based semantic segmentation yet, vehicles and pedestrians that remain still are classified as static objects.
As an additional demonstration of this dynamic object removal pipeline we have tested it against one of the sequences of the ERASOR dataset introduced in Lim et al. (2021). This dataset is specifically designed to capture situations with dynamic objects such as pedestrians or vehicles in motion. The same parameters were used than in the Paris LuCo dataset above, and results are shown in Figure 10. It can be verified how all traces of vehicles and pedestrians are correctly removed in the “static” layer.